Your grandfather placed $5,000 in a trust fund for you today. 10 years later the fund will be worth $8,000. What is the interest rate earned on this trust fund? (Annual compounding)
Answer #1

Present value = 5,000

Time = 10 years

Future value = 8000

FV = PV / (1+R)^N

(1+R)^N = 8000 / 5000

(1+R)^N = 1.6

(1+R)^10 = 1.6

1+R = 1.6^(1/10)

1+R = 1.048122

R = 0.048122

Hence, interest rate = 4.8122%